OpenStudy (anonymous):

Which decimal is equivalent to ? A. 8.12 B. 8.34 C. 8.75 D. 8.9

OpenStudy (danjs):

the copy paste didn't include everything....

OpenStudy (anonymous):

WHICH DECIMAL IS EQUIVALENT TO 8 9/12

OpenStudy (danjs):

9/12 = 3/4 = 0.75

OpenStudy (danjs):

8.75

OpenStudy (anonymous):

THANK YOU

OpenStudy (danjs):

YOU ARE WELCOME